# cryoSPHERE: Single-particle heterogeneous reconstruction from cryo EM

### v1

@@ 3.4 Decoder architecture
The decoder describes the distribution of the images given the latent variables. The latent variables include:
@@
    The global parameters $\{\mu_{m},\sigma_{m},\alpha_{m}\}_{m=1,\dots,N_{\textrm{segm}}}$ of the GMM describing the segment decomposition.
Given these latent variables and a base structure $S_{0}$, we parameterize the decoder $f_{\theta}$ in three steps.
First, a neural network with parameters $\theta$ maps $z_{i}\in\mathbb{R}^{L}$ to one 3D translation per segment $\{\vec{t}_{m}\}_{m=1,\dots,N_{\textrm{segm}}}$ and one rotation per segment. The rotation is represented by a unit quaternion $\{\vec{q}_{m}\}_{m=1,\dots,N_{\textrm{segm}}}$ (Vicci, 2001), which can be equivalently described as a set of axes of rotation $\{\vec{\phi}_{m}\}_{m=1,\dots,N_{\textrm{segm}}}$ and corresponding angles of rotations $\{\delta_{m}\}_{m=1,\dots,N_{\textrm{segm}}}$.
Second, given the parameters of the GMM, we can compute the matrix $G$. Finally, for each residue $i$ of $S_{0}$, we update the coordinates of all its atoms $\{a_{ik}\}_{k=1,\dots,A_{i}}$:
@@
    Second, it is translated according to:
    $a_{ik}^{\prime\prime}=a_{ik}^{\prime}+\sum_{j=m}^{N}G_{im}\vec{t}_{m}.$ (7)
This way, the rotation angles and translations for a residue incorporate contributions from all segments, proportionally on how much they belong to the segments. If condition (5) is met, a roughly rigid motion for each segment can be expected.
@@ 3.5 Encoder and priors
Figure 5: MD dataset. Left: Histograms of the distances of the two upper domains. The true distances are in green. The recovered distances are in blue. Right: Predicted against true distances in Ångström. The black line represent $x=y$.
We follow the classical variational-autoencoder framework. The distribution $q_{\psi}(y|I^{\star})$ is given by a normal distribution $\mathcal{N}(\mu(I^{\star}),\mathrm{diag}(\sigma^{2}(I^{\star})))$ where $\mu\in\mathbb{R}^{L}$ and $\sigma\in\mathbb{R}_{+}^{L}$ are generated by a neural network with parameters $\psi$, taking an image $I^{\star}$ as input. Additionaly, the approximate posterior distribution on the parameters of the GMM is chosen to be Gaussian and independent of the input image:
@@
where $\{\nu_{\mu_{m}},\beta_{\mu_{m}},\nu_{\sigma_{m}},\beta_{\sigma_{m}},\nu_{\alpha_{m}},\beta_{\alpha_{m}}\}_{m=1,\dots,N_{\textrm{segm}}}$ are parameters that are directly optimized and in practice we feed $\sigma_{m}$ to an ELU+1 layer in order to avoid a negative or null standard deviation.
@@
Finally, we assign standard Gaussian priors to both the local latent variable $z_{i}$ $z_{i}\sim\mathcal{N}(0,I_{L}),$ and the global GMM parameters $\{\mu_{m},\sigma_{m},\alpha_{m},\}_{m=1,\dots,N_{\textrm{segm}}}$.
Note that applying the reparameterization trick (Kingma & Welling, 2014) is straightforward for a Gaussian distribution. Calculating the KL-divergence between two Gaussian distributions as in Equation (4), is also straightforward.
@@ 4 Experiments
In this section, we test cryoSPHERE on different datasets and compare the results to cryoDRGN (Zhong et al., 2020). CryoDRGN is a state-of-the-art method for continuous heterogeneous reconstruction, in which the refinement occurs at the level of electron densities. To synthesize data for all experiments, we applied CTF corruption to the images and add Gaussian noise to achieve $\mathrm{SNR}\approx 0.1$. See Appendix A for a detailed description of the CTF. Furthermore, we make the assumption that we know the exact poses. However, since we use a structure $S_{0}$, which is different from the structures used to generate the datasets, these poses can only be an approximation. This will not be the case of cryoDRGN, for which these poses will indeed be exact, as this method does not use a base structure. Consequently, in this context, the comparison may introduce a bias in favour of cryoDRGN.
4.1 Toy dataset: a continuous motion.
For an initial assessment of cryoSPHERE, we examine a simulated dataset based on a bacterial phytochrome with UniProt (The UniProt Consortium, 2021) entry Q9RZA4. This protein is a dimer, with 755 residues per chain. To simulate a continuous heterogeneous set of images, we predicted the structure with AlphaFold multimer (Evans et al., 2021) and divided the protein into two domains. The first one consists of the full chain A and 598 residues of chain B, 1353 residues in total. The second domain comprises the last 157 residues of the chain B, see Figure 3. We then sample $10^{4}$ rotation angles from a two-components Gaussian mixture. Each conformation is generated by rotating the second domain according to the axis $(0,1,0)$ and one of the sampled rotation angles. This yields $10^{4}$ different conformations. Finally, we uniformly sample $15$ different poses for each structure, yielding a total of $150$k images. Appendix A.1 provides more details, and Figure 3 shows examples of structures.
Moreover, Figure 3 shows that most of the predicted angles of rotation for the fourth segment are in excellent agreement with the ground truth structural changes. However, approximately $12$ of the dataset is predicted as belonging to the wrong mode. Figure 8 in the appendix shows that the norms of the predicted translations for both segments are negligible, also indicating that cryoSPHERE has recovered the ground truth. The same figure shows that we recover the right axis of rotation. In Appendix A.1, we plot the predicted angles against the mean of the latent variable distributions output by the encoder. We observe that cryoSPHERE has recovered the circular motion: the greater the latent mean, the lower the predicted angle, with a quick transition from one mode to the other. This indicates that cryoSPHERE is able to recover the structure of single-particles from cryo EM data.
4.2 A discrete conformation dataset: the SARS-CoV-2 spike protein
We now test cryoSPHERE on a dataset comprising two discrete conformations of the SARS-CoV-2 spike protein: a close conformation with PDB entry 6ZP5 and an open conformation with PDB entry 6ZP7, see Melero et al. (2020). We generate $150$k images of size $N_{\textrm{pix}}=300$ with uniformly sampled poses and neglected translations, with seventy percent corresponding to the close conformation and thirty percent to the open one.
Prior to running cryoSPHERE, we obtain a base structure by first running the multimer version of AlphaFold (Evans et al., 2021) on the amino acid sequence of the spike protein, selecting the best (according to AlphaFold) structures. Second, we utilize a custom backprojection algorithm provided by (Zhong et al., 2020) to get a homogeneous reconstruction. This yields a volume with low resolution in the mobile parts but decent resolution in the less mobile ones. To reduce a potential bias introduced by using an AlphaFold structure, we subsequently fit this structure into the volume, initially performing rigid-body fitting in ChimeraX (Meng et al., 2023) and then using ISOLDE (Croll, 2018). Subsequently, we run cryoSPHERE with this fitted structure as the base structure for 48 hours on a single GPU, totalling 174 epochs.
We define the two conformations in terms of the angle between two domains, see Figure 4. The same figure demonstrates that we recover the angles within roughly $1$ degree. After clustering the structures predicted by cryoSPHERE based on their angle defined in Figure 4, only $4$ of them get predicted to be in the wrong conformation. Additionally, by turning the structures predicted by cryoSPHERE into volumes using our operator (1), we observe excellent agreement with the ground truth volumes, as shown by the Fourier shell correlation (FSC) curves (Klaholz, 2015). This illustrates that cryoSPHERE can effectively recover discrete conformational disorder.
@@ 4.2 A discrete conformation dataset: the SARS-CoV-2 spike protein
We now test cryoSPHERE with a more difficult dataset. This time, we simulate a continuous motion of a bacterial phytochrome, with PDB entry 4Q0J (Burgie et al., 2014). This dimer has 503 residues per chain. We define two upper domains corresponding to the residues 321-502 from both chain A and B. Then, we run a Molecular Dynamics (MD) simulation to simulate a dissociation of these two domains, as depicted in Figure 9 in Appendix A.2. We take $10^{4}$ structures along the trajectory and sample $15$ poses uniformly for each structure. Refer to Appendix A.2 for more details.
Figure 6: MD dataset. Left: Recovered segments on the fitted AlphaFold structure. The colors denotes different contiguous domains: if two segments share the same color, it does not mean their residues belong to each others. Middle left to middle right: cryoSPHERE predicted structures in green, corresponding ground truth in blue. Right: Fourier shell correlation curves. The mean across structures for cryoSPHERE is in plain blue. Dotted blue represents the mean $\pm$ one standard deviation. The red, plain line represents the mean for cryoDRGN across volumes, with the dotted red line representing the mean $\pm$ one standard deviation.
Analogously to Section 4.2, we obtain an AlphaFold structure that we fit into a volume reconstructed by backprojection. Further details can be found in Appendix A.2. Finally, we run our method with the fitted AlphaFold structure as our base structure and $N_{\textrm{segm}}=20$.
Since cryoSPHERE outputs structures, we can compute the predicted distances between the above-mentioned domains. We plot the true and recovered histograms of the distances between the two domains in Figure 5. Additionally, we plot the predicted distances against the true distances in the same figure. These two plots show that the predicted structures are in excellent agreement with the ground truth structures. This is even true for the most dissociated structures, even though the majority of the structures are in a closed conformation. Figure 6 shows the recovered segments as well as a selection of predicted structures and the corresponding ground truths. Our model has learnt to separate the two mobile top domains from the fix bottom one.
For comparison, we train cryoDRGN for the same duration using the same single GPU. We use the default settings of cryoDRGN, as detailed in Appendix A.2.
We plot the mean of the FSC curves between the predicted volumes and the corresponding ground truth volumes in Figure 6, for both cryoSPHERE and cryoDRGN. This plot shows that we perform better than cryoDRGN at both the $0.5$ and $0.143$ cutoffs. There are two reasons behind this.
First, we fit our base structure into a volume reconstructed from a homogeneous reconstruction method. This aligns the non-moving parts of the protein with the right places, as provided by the dataset. Without this step, it is likely that some medium-scale part of the protein would have been misplaced, potentially reducing the performance of our algorithm at the $0.5$ cutoff; see Appendix A.2 for evidence of this. It is important to note, that this step significantly improves our results, and it comes at almost no extra cost since it relies on the backprojection algorithm.
Secondly, deforming a good AlphaFold structure offers finer details than cryoDRGN. This accounts for the better performance of cryoSPHERE over cryoDRGN at the $0.143$ cutoff. Refer to Appendix A.2 for examples of reconstructed structures and volumes.
Moreover, our method demonstrates robustness to a limited number of images. Even with only one pose per structure (totalling $10^{4}$ images), cryoSPHERE maintains similar performance, while cryoDRGN suffers a significant drop in the FSC curves. Further details can be found in Appendix A.2.

### v2

@@ 3.4 Decoder architecture
The decoder describes the distribution of the images given the latent variables, which include:
@@
    The global parameters $\{\mu_{m},\sigma_{m},\alpha_{m}\}_{m=1}^{N_{\textrm{segm}}}$ of the GMM describing the segment decomposition.
Given these latent variables and a base structure $S_{0}$, we parameterize the decoder $f_{\theta}$ in three steps. First, a neural network with parameters $\theta$ maps $z_{i}\in\mathbb{R}^{L}$ to a set of rigid body transformations, one for each segment $m=1,\dots,N_{\textrm{segm}}$. The transformation of segment $m$ is represented by a translation vector $\vec{t}_{m}$ and a unit quaternion $\vec{q}_{m}$ (Vicci, 2001), which can further be decomposed into an axis of rotation $\vec{\phi}_{m}$ and rotation angle $\delta_{m}$. Second, given the parameters of the GMM, we compute the matrix $G$. Finally, for each residue $i$ of $S_{0}$, we update the coordinates of all its atoms $\{a_{ik}\}_{k=1}^{A_{i}}$:
@@
    Second, it is translated according to: $a_{ik}^{\prime\prime}=a_{ik}^{\prime}+\sum_{j=m}^{N}G_{im}\vec{t}_{m}.$
This way, the transformation for a residue incorporate contributions from all segments, proportionally on how much they belong to the segments. If condition (5) is met, a roughly rigid motion for each segment can be expected.
@@ 3.5 Encoder and priors
Figure 3: MD dataset SNR 0.001. Left: Histograms of the distances of the two upper domains. The true distances are in green. The recovered distances are in blue. Right: Predicted against true distances in Ångström. The black line represent $x=y$.The correlation between the predicted and true distances is 0.73. For the same plot for cryoStar, see Appendix B.2 of the supplementary file.
We follow the classical VAE framework. The distribution $q_{\psi}(y|I^{\star})$ is given by a normal distribution $\mathcal{N}(\mu(I^{\star}),\mathrm{diag}(\sigma^{2}(I^{\star})))$ where $\mu\in\mathbb{R}^{L}$ and $\sigma\in\mathbb{R}_{+}^{L}$ are generated by a neural network with parameters $\psi$, taking an image $I^{\star}$ as input. Additionally, the approximate posterior distribution on the parameters of the GMM is chosen to be Gaussian and independent of the input image:
@@
where $\{\nu_{\mu_{m}},\beta_{\mu_{m}},\nu_{\sigma_{m}},\beta_{\sigma_{m}},\nu_{\alpha_{m}},\beta_{\alpha_{m}}\}_{m=1}^{N_{\textrm{segm}}}$ are parameters that are directly optimized. In practice we use ELU+1 layers for $\sigma_{m}$ to avoid negative or null standard deviation.
@@
Finally, we assign standard Gaussian priors to both the local latent variable $z_{i}\sim\mathcal{N}(0,I_{L}),$ and the global GMM parameters $\{\mu_{m},\sigma_{m},\alpha_{m},\}_{m=1}^{N_{\textrm{segm}}}$. This reparameterization (Kingma & Welling, 2014) is straightforward for a Gaussian distribution. Calculating the KL-divergence between two Gaussian distributions as in equation 4, is also straightforward.
3.6 Loss
Since the images may be preprocessed in unknown ways before running cryoSPHERE, we use a correlation loss between predicted and ground truth image instead of a mean squared error loss, similar to (Li et al., 2023):
$\mathcal{L}_{\mathrm{corr}}=\dfrac{-I_{i}^{\star}\cdot I(R_{i},t_{i},f_{\theta}(S_{0},z))}{||I_{i}^{\star}||\times||I(R_{i},t_{i},f_{\theta}(S_{0},z))||}$ (7)
where $\cdot$ denotes the dot product. The total loss to minimize writes:
$\mathcal{L}(I,I^{\star})=\mathcal{L}_{\mathrm{corr}}+\mathrm{D}_{\mathrm{KL}}(q_{\psi}(z|I^{\star})||p(z))$ (8)
In our experience, it is unnecessary to add any regularization term to the correlation and KL divergence losses, except for datasets featuring a very high degree of heterogeneity. In that case, we offer the option of adding a continuity loss to avoid breaking the protein and a clashing loss to avoid clashing residues, as it is done in (Rosenbaum et al., 2021; Li et al., 2023; Jumper et al., 2021). We describe these losses in Appendix A.1 of the supplementary file.
4 Related works
Two of the most popular methods for cryo-EM reconstruction, which are not based on deep learning, are RELION (Scheres, 2012) and cryoSPARC (Punjani et al., 2017). Both methods perform volume reconstruction, hypothesize that $k$ conformations are present in the dataset and perform maximum a posteriori estimation over the $k$ density maps, thus performing discrete heterogeneous reconstruction. Both of these algorithms operate in Fourier space using an expectation-maximization algorithm Dempster et al. (1977) and are non-amortized: the poses are refined for each image. Other approaches perform continuous heterogeneous reconstruction. For example, 3DVA (Punjani & Fleet, 2021b) uses a probabilistic principal component analysis model to learn a latent space.
Another class of methods involve deep learning and typically performs continuous heterogeneous reconstruction using a VAE architecture. Of those that attempt to reconstruct a density map, cryoDRGN (Zhong et al., 2020; 2021a) and CryoAI (Levy et al., 2022) use a VAE acting on Fourier space to learn a latent space and a mapping that associates a 3D density map with each latent variable. They perform non-amortized and amortized inference over the poses, respectively. Other methods are defined in the image space, e.g. 3DFlex (Punjani & Fleet, 2021a) and cryoPoseNet (Nashed et al., 2021). They both perform non-amortized inference over the poses. These methods either learn, for a given image $I_{i}$, $\{V_{i}(x_{k})\}$ the values at a set of $N_{\textrm{pix}}^{3}$ fixed 3D coordinates $\{x_{k}\}$, representing the volume on a grid (explicit parameterization), or they learn an actual function $\hat{V}_{i}:\mathbb{R}^{3}\rightarrow\mathbb{R}$ in the form of a neural network that can be queried at chosen coordinates (implicit parameterization). These volume-based methods cannot use external structural restraints or force fields as additional information. This limits their applicability to low SNR data sets, which are frequent in protein cryo EM.
Other deep learning methods attempt to directly reconstruct structures instead of volumes and share a common process: starting from a plausible base structure, obtained with e.g. AlphaFold (Jumper et al., 2021), for each image, they move each residue of the base structure to fit the conformation present in that specific image. These methods differ on how they parameterize the structure and in the prior they impose on the deformed structure or the motion of the residues. For example AtomVAE (Rosenbaum et al., 2021) considers only residues and penalizes the distances between two subsequent residues that deviate too much from an expected value. CryoFold (Zhong et al., 2021b) considers the residue centers and their side-chain and also imposes a loss on the distances between subsequent residues and the distances between the residue centers and their side-chain. Unfortunately, due to the high level of noise and the fact that we observe only projections of the structures, these ”per-residue transformation” methods tend to be stuck in local minima, yielding unrealistic conformations unless the base structure is taken from the distribution of conformations present in the images (Zhong et al., 2021b), limiting their applicability on real datasets. Even though AtomVAE (Rosenbaum et al., 2021) could roughly approximate the distribution of states of the protein, it was not able to recover the conformation given a specific image.
To reduce the bias that the base structure brings, DynaMight (Schwab et al., 2023) fits pseudo-atoms in a consensus map with a neural network directly. Similar to our work, several other methods constrain the atomic model to rigid body motions. For example e2gmm (Chen & Ludtke, 2021; Chen et al., 2023) deform a nominal structure $S_{0}$ based on how much its residues are close to a learnt representation $S_{\mathrm{small}}$ of $S_{0}$. This is similar to our GMM, except that their takes place in $\mathbb{R}^{3}$ and is not used to perform rigid body motion. Instead, they ask the user to define the segmentation in a later step. This is in contrast to cryoSPHERE, which learns the motion and the segmentation concurrently. Using DynaMight (Schwab et al., 2024), Chen et al. (2024) developed a focused refinement on patches of the GMM representation of the protein. These patches are learnt using $k$-means on the location of residues and do not depend on the different conformations of the data set. This in contrast to cryoSPHERE where the learning of the segments of the protein is tightly linked to the change of conformation. Concurrently to our work, Li et al. (2023) developed cryoStar which learns to translate each residue independently using a variational auto-encoder. They enforce the local rigidity of the motion of the protein by imposing a similarity loss between the base structure and the deformed structure as well as a clash loss. The interested reader can see Donnat et al. (2022) for an in-depth review of deep learning methods for cryo-EM reconstruction.
The reconstruction methods relying on an atomic model, such as cryoStar, DynaMight or cryoSPHERE offer the possibility to the user to provide prior information via this atomic model. They also offer the possibility of deforming the protein according to chemical force fields. This is not the case of the methods performing volume reconstruction without such an atomic model.
Figure 4: MD dataset. Left: cryoSPHERE Recovered segments. The colors denotes different contiguous domains. Middle and right: mean FSC comparison +/- one standard deviation, for cryoSphere and cryoDRGN and cryoStar. For a comparison between cryoStar and cryoDRGN, see Appendix B.2 in the supplementary file.
@@ 5 Experiments
In this section, we test cryoSPHERE on a set of synthetic [^2] See Appendix B of the supplementary file for details on how we created the synthetic datasets. and real datasets with varying level of noise and compare the results to cryoDRGN (Zhong et al., 2020) and cryoStar (Li et al., 2023). CryoDRGN is a state-of-the-art method for continuous heterogeneous reconstruction, in which the refinement occurs at the level of electron densities, while cryoStar is a structural method similar to ours. To our knowledge, the code for AtomVAE and CryoFold is not available and non-trivial to reimplement. For this reason we focus our comparison on the aforementioned methods, which have furthermore reported state-of-the-art performance. In Appendix B.1, we demonstrate that cryoSPHERE is able to recover the exact ground truth when it exists. We also discuss its performances with varying SNR and $N_{\textrm{segm}}$ and show how to debias cryoSPHERE results using DRGN-AI or cryoStar volume method in Appendix B.2. Finally, Appendix B.5 compares the computational costs of cryoSPHERE and cryoStar.
@@
As a more difficult test case we simulate a continuous motion of a bacterial phytochrome, with PDB entry 4Q0J (Burgie et al., 2014). The trajectory starts at the closed conformation of Figure 11 and ends at the most open conformation on the same figure. It corresponds to a dissociation of the two top parts of the protein. This dataset has a very low SNR of $0.001$. Our base structure is obtained by AlphaFold and is subsequently fitted into a homogeneous reconstruction given by the backprojection algorithm. We train cryoSPHERE with $N_{\textrm{segm}}=25$, cryoStar, and cryoDRGN for 24 hours each, using the same single GPU. We get one predicted structure per image for cryoSPHERE and cryoStar, that we turn into volumes using (1), and one predicted volume per image for cryoDRGN. See Appendix B.2 in the supplementary file for details and comparison with different values of $N_{\textrm{segm}}$. Note that since both cryoSPHERE and cryoStar use a nominal structure, we fit the structure we obtained through AlphaFold in the consensus reconstruction obtained by backprojection and use that exact same structure as the nominal one for both methods.
Figure 3 shows the predicted distance between the two upper parts of the protein being dissociated, against the ground truth distance for each image. In spite of the very low SNR, cryoSPHERE roughly recovers the right distribution of distances. More importantly, the correlation between the predicted distance and ground truth distance is $0.74$, showing that cryoSPHERE is able to recover the correct conformation given an image. This is in stark contrast with Rosenbaum et al. (2021) who could not recover the conformation conditionally on an image. In addition, our model has learnt to separate the two mobile top domains from the fix bottom one, as shown by the segment decomposition in Figure 4. Appendix B.2 in the supplementary file shows the same figures for cryoStar.
We plot the mean of the FSC curves between the predicted volumes and the corresponding ground truth volumes in Figure 4, for cryoSPHERE, cryoDRGN and cryoStar. CryoSPHERE performs better than both cryoDRGN and cryoStar at both the $0.5$ and $0.143$ cutoffs. We attribute this to three key properties. Firstly, we fit our base structure into a consensus reconstruction. This step corrects the position of the medium-scale elements of the base structure that could have been misplaced, boosting the FSC of cryoSPHERE at the $0.5$ cutoff. Secondly, acting directly on the structure level offers a finer resolution than cryoDRGN given the level of noise. Figure 32 shows that cryoDRGN underestimated the opening of the protein and sometimes gives very noisy volumes. That explain why we outperform cryoDRGN at the $0.143$ cutoff. Finally, cryoSPHERE is rigidly moving larger segments of the protein. This provide a better resistance to high levels of noise and overfitting compared to moving each residue individually like cryoStar does, providing a possible explanation to the improvement compared to cryoStar at the $0.143$ cutoff.
5.2 EMPIAR 10180
[…]
Figure 5 shows a set of ten structures taken evenly along the first principal component of the latent space. To interrogate if these structures contain bias from the structural constraints, we perform a volume reconstruction step similar to cryoStar Phase II, see Figure 5.
Traversing the first principal component shows that the Sf3b domain gets incurvated down while the helicase move closer to the foot of the protein. This is in line with the literature (Li et al., 2023; Plaschka et al., 2017). The motion of the protein also brings the alpha helix of the Spp381 domain closer to the foot, as corroborated by Li et al. (2023). Comparison between the recovered structures and volumes (Figure 50) shows similar movements, indicating a small amount of bias from the structural constraints. In addition, the absence of density corresponding to the U2 domain in the volume indicates that it there is compositional heterogeneity that cryoSPHERE could not detect, see Figure 5. We provide a movie of the motion and more structures and volumes in appendix B.3 in the supplementary file.
Figure 5: EMPIAR10180. Left and middle left: different views of the structures corresponding to the red dots of Figure 48. The motion goes from red (left in the first principal component) to white to blue (right of the principal component). Only the $C_{\alpha}$ atoms are shown. Right and middle right: different views of two volumes recovered by training DRGN-AI on the latent space of cryoSPHERE. The U2 domain disappears on the volume because of a compositional heterogeneity.
5.3 EMPIAR-12093
We now tackle the recently published EMPIAR-12093 (Bódizs et al., 2024). This dataset comprises two sets of images: one non-activated (Pfr) and one activated (Pr). These dataset are very challenging because of the high level of noise and heterogeneity of the protein, especially in the Pfr dataset. Traditional methods like cryoSparc (Punjani et al., 2017) or cryoDRGN (Zhong et al., 2020; 2021a) fail at reconstructing the upper part of the protein, see Bódizs et al. (2024) and Appendix B.4.
Figure 6: EMPIAR12093. Left of the black line: Ten structures sampled along PC1, for cryoSPHERE and cryoStar. Right of the black line: examples of volumes reconstructed by training the cryoStar volume method on the latent space of cryoSPHERE for debiasing. The blue and red volumes correspond to the first and last volumes along PC1. Top row corresponds to Pr; bottom row to Pfr.
Figure 6 shows principal component 1 traversal for cryoStar and cryoSPHERE. For Pr, both methods are in strong agreement and reveal a rotation of the upper domain around its axis, while the lower part remains stationary. This aligns with previous studies (Wahlgren et al., ; Malla et al., 2024)).
Figure 7: EMPIAR12093. Distribution of the number of clashes for 2000 randomly chosen structures Pfr dataset, for cryoSPHERE and cryoStar. Two non contiguous residues are said to be clashing if their distance is less than 4 Å.
The Pfr dataset showcases an even lower SNR and more dynamical protein: the protein opens up completely. From consensus reconstructions alone, one could suspect that the upper domains are cut off in the sample preparation procedure. However, the protein is complete in Pr (light-activated) structure and the photocycle is reversible,(Takala et al., 2014) suggesting that this is not the case and that strong conformational heterogeneity that is at play.
For Pfr cryoStar is unable to produce physically plausible results: the top part of the protein appears disordered and shows a random motion. In addition, cryoStar does not recover the ”scissoring” motion of protein, which is thought to be active (Bódizs et al., 2024). On the contrary, cryoSPHERE gives a high level of motion in a structured manner and recovers the ”scissors” opening of the protein. Without any clashes (Fig. 7). (Bódizs et al., 2024).
Analysis of the dataset on phytochromes illustrates the scope and limitations of the different methods. Pure image-based methods (i.e. cryo DRGN) already fail on the Pr state with its intermediate disorder, while cryoSTAR and cryoSPHERE succeed in obtaining reasonable reconstructions (Figure 6). For the Pfr state it becomes evident that cryoSTAR struggles with the high noise and large motions encoded in the dataset. Its deformation-based approach results in unphysical motions along the first principal component, often leading to structural clashes. In contrast, cryoSPHERE handles the noise effectively, producing physically plausible large-scale motions in both the upper and lower domains, see Figure 7, the supplementary movies and B.4. We assign this superior performance to the higher degree of structural constraints that are used in cryoSPHERE compared to cryoSTAR.
We also performed debiasing of cryoSPHERE with a volume method and show examples of reconstructed volumes in Figure 6. For Pr, recovered densities are visible for the entire protein and confirm the dynamics of the upper domains, confirming the absence of compositional heterogeneity and a minimum of bias due to structural constraints. However, for Pfr, meaningful density of the upper (dynamic) part of the protein cannot be recovered, because the signal level in the averaged density is too low. Thus, for this most dynamic protein case, volume-based debiasing is not possible, despite the fact that the structure based cryoSPHERE finds solutions that fit the data set.